Give the center of the circle with equation x^2+2x+y^2-10y+22=0. Thank you!
Answer by jake_6233(74)

Give the center of the circle with equation x^2+2x+y^2-10y+22=0
You need to complete the square twice to solve this problem
Recall (x+b/2)^2 +(y+b/2)^2 then remember to add (b/2)^2 to the right side so
(x+1)^2 + (y-5)^2 = -22 +1 + 25
(x+1)^2 +(y-5)^2 = 4 (this is the standard form of a circle)
(x-h)^2 +(y-c)^2 = r^2
so the center is (-1,5) with a radius of 2